−y−10=6x
5x+y=−10

Answer by Theo(13342)

your equations are:
-y - 10 = 6x
5x + y = -10

rearrange the first equation so that x and y are on one side of the equation and the constant term is on the other side.

if you add y to both sides of that equation, you get -10 = 6x + y
this is the same as 6x + y = -10

your two equations becomes:

6x + y = -10
5x + y = -10

subtract the second equation from the first to get:

x = 0

in the first equation, solve for y to get y = -10

when x = 0 and y = -10, you get:

-y - 10 = 6x which becomes 10 - 10 = 0 which is true.
this is the original first equation.

5x + y = -10 which becomes 0 - 10 = -10 which becomes -10 = -10 which is true.
this is the original second equation.

the values of x and y that satisfy both original equations are x = 0 and y = -10.